OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x3 + 4x2 + x – 6

OpenStudy (anonymous):

this one is easier

OpenStudy (anonymous):

only possible rational zeros are \[\pm1,\pm2,\pm3,\pm6\]

OpenStudy (anonymous):

you can see that one works with your eyeballs. so factor it out to get the other zeros

OpenStudy (anonymous):

\[x^3-4x^2+x-6=(x-1)(x^2+5x+6)=(x-1)(x+2)(x+3)\]